OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2).

OpenStudy (anonymous):

You can simply plug that point into the left side of the equation, and obtain \(3(4)-(-2)=?\), then solve to obtain the line \(3x-y=14.\)

OpenStudy (anonymous):

Parallel lines have the same slope. First solve for y here to get the slope. y = 3x + 3 THe slope is 3. y = 3x + b Plug in the coordinate. -2 = 3(4) + b -2 = 12 + b b = -14 y = 3x - 14

OpenStudy (mertsj):